OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6).

OpenStudy (anonymous):

y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (luigi0210):

Parallel, use the same slope: \(m=-3\) Use the point slope formula: \(\LARGE y-y_{1}=m(x-x_{1})\)

OpenStudy (anonymous):

Where \[(x_1,y_1) \] are your coordinates, and x and y are variables, solve for y.

OpenStudy (anonymous):

Aline parallel to another line must have the same slope. The only difference can therefore be the y intercept of the line. The slope of original line = -3 The equation of the parallel line is y = -3x +C where c is the intercept of the parallel line. To find C we have a clue. The new line must pass through (-1, 6) So the equation when we substitute for x and y becomes, 6 = -3 * -1 + C C = 6 -3 = 2 So the parallel line equation is y = -3x + 2

OpenStudy (anonymous):

Would it be B?